Jill has $3.50 in nickels and dimes. If she has 50 coins, how many of each type of coins does she have? I need to know an X and a Y. THANK YOU
Answer by Earlsdon(6294)

Let x = the number of dimes ($0.10) and y = the number of nickels ($0.05).
x + y = 50 The total number of coins. Rewrite this as: x = 50-y and substitute for x in the second equation.
($0.10)x + ($0.05)y = $3.50
($0.10)(50-y) + ($0.05)y = $3.50 Simplify and solve for y.
5 - 0.1y + 0.05y = 3.5
5 - 0.05y = 3.5 Subtract 5 from both sides.
-0.05y = -1.5 Divide both sides by -0.05
y = 30 This is the number of nickels.
x = 50-y
x = 50-30 = 20 This is the number of dimes.